# The Unbounded Denominators Conjecture

### v1

@@ 5.5. The region $F^{N}_{N}\sim 1$ and $F_{N}\sim\infty$
For $N\in\mathbf{N}$ and $r\in(0,1)$, we have
$\sup_{|z|=r}\log|F_{N}|\ll\frac{N}{1-r},$
where the implicit coefficient is absolute.
@@
By Lemma 5.3.1, the set $\{z\in D(0,1)\,:\,|F_{N}(z)|>e^{M}+1\}$ is contained in the rotation under $s$ of the set $S(M,N):=\{z\in D(0,1)\,:\,|F_{N}^{N}(z)-1|<e^{-MN}\}$. Thus to prove the lemma, it suffices to show that for $\frac{N}{1-r}\gg 1$, if we set $M=\frac{N}{1-r}$, then $S(M,N)$ is contained in $\{z\in D(0,1)\,:\,|z|>r\}$. (When $\frac{N}{1-r}\leq C$ for some large constant absolute $C$, then $\sup_{|z|=r}\log|F_{N}|\leq\sup_{|z|=1-1/C}\log|F_{N}|$, which is a finite number and there are only finitely many $N$ to consider; thus the lemma follows trivially in this case.)
It follows from Lemma 5.5.1 that
$S(M,N)\subset\bigcup_{\gamma\in\Gamma_{N}}\gamma\mathcal{H}_{D},\quad\textrm{where }D=\frac{N^{2}M}{\pi^{2}}.$
Thus we only need to show that the diameters $E(\gamma,D)\leq 1-r$ of all $\gamma\mathcal{H}_{D}$.
By Shimizu’s Lemma (see, for example, [EGM98, Theorem 3.1]), we have  $2|c|\cot(\pi/2N)\geq 1$ for all $\widetilde{\gamma}=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\widetilde{\Gamma}_{N}$. Thus by Lemma 5.5.5,
@@
where the implicit constant is absolute. Thus we have $E(\gamma,D)\leq 1-r$ once $N^{-1}$ times the implicit constant is less than $1$. ∎
@@ 6. Nevanlinna theory and uniform mean growth near the boundary
For our application of Theorem 2.0.2, we prove in this section the following uniform growth bound. Throughout this section, we assume as we may that $N\geq 2$. Then the analytic map $F_{N}:D(0,1)\to\mathbf{P}^{1}$ omits the $N+1\geq 3$ values $\mu_{N}\cup\{\infty\}$. In such a situation, we seek to exploit whatever growth constraints are imposed on the map by Nevanlinna’s value distribution theory. A theorem of Tsuji [Tsu52, Theorem 11] gives the general asymptotic
@@
for any universal covering map $F:D(0,1)\to\mathbf{C}\smallsetminus\{a_{1},\ldots,a_{N}\}$ (see also the discussion in Nevanlinna [Nev70, page 272]), however this is only asymptotically in $r\to 1^{-}$ for given punctures $\{a_{i}\}$ whereas we need a uniformity in both $r$ and $N$. It is at the point (6.1.2) of the explicit partial fraction coefficients that our argument below makes a critical use of the special feature of the target set $\mu_{N}\cup\{\infty\}$ of omitted values.
@@
This quantity is known in Nevanlinna theory as the mean proximity function at $\infty$
@@
The Nevanlinna characteristic
$T(r,f):=m(r,f)+N(r,f)$
satisfies for every $a\in\mathbf{C}$ the relation
@@
6.1. Reduction to a logarithmic derivative
For $f:D(0,1)\to\mathbf{C}$ holomorphic as opposed to meromorphic, the polar divisor is empty and $N(r,f)=0$, so in that case $m(r,f)=T(r,f)$. Since by design $F_{N}^{N}-1$ is a unit in the ring of holomorphic functions on $D(0,1)$, our requisite bound (6.0.2) rewrites in Nevanlinna notation into
(6.1.1) $T(r,p\circ F_{N})\ll\log{\frac{N}{1-r}},\quad\textrm{for each of }p(x)\in\big\{x^{N}/(x^{N}-1),\,1/(x^{N}-1),\,x^{N}\big\},$
and Lemma 6.0.3 using $x^{N}/(x^{N}-1)=1+1/(x^{N}-1)$ and (6.0.6) shows that the later three cases for $p(x)$ are equivalent to one another. See the explicit inequality (6.1.5) below for one of these implications. We will prove Theorem 6.0.1 in the form $T(r,F_{N}^{N})\ll\log{\frac{N}{1-r}}$ but pivoting around the choice
(6.1.3) $f:=1-F_{N}^{N}\quad:\quad D(0,1)\to\mathbf{C}^{\times}$
is related to $p\circ F_{N}=F_{N}^{N}/(F_{N}^{N}-1)$ by
(6.1.4) $p\circ F_{N}=\frac{F_{N}}{NF_{N}^{\prime}}\frac{f^{\prime}}{f}.$
We furthermore have, since $F_{N}(0)=0$ and $F_{N}^{N}-1$ is a unit in the ring of holomorphic functions on $D(0,1)$:
$=T\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}=T(r,F_{N}^{N}-1)-\log{2}$
(6.1.5) $\geq T(r,F_{N}^{N})-2\log{2}=N\,T(r,F_{N})-\log{4}.$
Our proof of Theorem 6.0.1 combines Lemma 5.5.1 with the centerpiece of R. Nevanlinna’s original analytic proof — based on the lemma of the logarithmic derivative — of his second main theorem of value distribution theory. To bound the logarithmic derivative terms in (6.1.4), it is sufficient to cite [Nev70, Lemma IV.3.1 on page 244] or [Hay64, Lemma 2.3 on page 36]. For convenience to the reader, and particularly since the argument simplifies considerably in the case that we need of a functional unit (a nowhere vanishing holomorphic function), we include our own self-contained treatment of a basic explicit case of the lemma on the logarithmic derivative.
@@ 6.1. Reduction to a logarithmic derivative
The case of arbitrary meromorphic functions $g:\overline{D(0,R)}\to\mathbf{P}^{1}$ is handled similarly by a differentiation in the general Poisson–Jensen formula, but with rather more work to estimate the finite sum over the zeros and poles of $g$. In this way, by using a technique due to Kolokolnikov for handling the sum over the zeros and poles, Goldberg and Grinshtein [GG76] obtained the general bound
$m\Big(r,\frac{g^{\prime}}{g}\Big)<\log^{+}{\Big\{\frac{T(R,g)}{r}\frac{R}{R-r}\Big\}}+5.8501,\quad\textrm{for }g(0)=1,$
@@
On applying Lemma 6.1.6 to the nowhere vanishing holomorphic function (6.1.3) and the outer radius choice
R:=1(1r)/2=(1+r)/2,R:=1-(1-r)/2=(1+r)/2,
and using (cf. [BG06, Corollary 13.2.14]) that $m(r,f^{\prime}/f)=T(r,f^{\prime}/f)$ is a monotone increasing function of $r$, we find the mean growth bound
@@
(6.2.1) $\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log^{+}{\frac{N}{1-r}},$
where in the last step we have estimated a mean proximity function trivially by a supremum function.
We shall handle the term $F_{N}/F_{N}^{\prime}$ in (6.1.4) by employing Lemma 6.1.6 to the functional unit $g=1-F_{N}$, and the following standard chain of implications based on Jensen’s formula in the reduction of the second main theorem to the lemma of the logarithmic derivative (see, for example, [Hay64, pages 33–34]), beginning with (6.0.5) for the function $F_{N}^{\prime}/F_{N}$, and using that our function $F_{N}$ is holomorphic on the disc $D(0,1)$ with $F_{N}(0)=0$ and $F_{N}^{\prime}(0)\neq 0$:
@@
$=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big).$
Here for the last equality we recall that $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ is an étale analytic mapping, hence the derivative $F_{N}^{\prime}$ is nowhere vanishing and the ramification term $N_{\mathrm{ram}}(r,F_{N})=N(r,1/F_{N}^{\prime})=0$ is actually zero. (In any event one could drop a ramification term $N_{\mathrm{ram}}=N_{1}\geq 0$ by positivity, here or in [Hay64, Theorem 2.1]. See also the discussion in Remark 6.2.5 below.)
@@
(6.2.2) $\leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big),$
upon again using Lemma 6.1.6 with R:=(1r)/2R:=(1-r)/2 but now for the functional unit $g:=1-F_{N}$, and a similar argument as in (6.2.1).
At this point the key identity (6.1.4) allows to combine the estimates (6.2.1) and (6.2.2), giving at the uniform bound
@@
outside of an exceptional set of radii $r\in E\subset[0,\infty)$ of finite Lebesgue measure: $m(E)<\infty$. Here $N_{\mathrm{ram}}(r,g)=N(r,1/g^{\prime})$ is a ramification term, which is always nonnegative and vanishes if the map $g$ is étale. This is Nevanlinna’s quantitative strengthening of Picard’s theorem on at most one omitted value for a nonconstant entire function, for if each of $a_{1},\ldots,a_{N}$ is omitted then all counting terms $N(r,a_{i})=0$ vanish on the right-hand side of (6.2.6), leading if $N\geq 2$ to an $O(1)$ upper bound on the growth $T(r,g)$ of $g$. The idea is that we similarly have a holomorphic map $F_{N}$ omitting the $N$ values $a_{h}=\exp(2\pi ih/N)$, except $F_{N}$ is on a disc rather than the entire plane, and that (6.2.6) largely extends as a growth bound for holomorphic maps on a disc. For such completely quantitative results we refer the reader to Hinkkanen [Hin92, Theorem 3] or Cherry–Ye [CY01, Theorem 4.2.1 or Theorem 2.8.6]. We cannot directly apply these general theorems in their verbatim forms as they only lead to a bound of the form $m(r,F_{N})\ll\log{\frac{N}{1-r}}$ in place of the required $m(r,F_{N})\ll\frac{1}{N}\log{\frac{N}{1-r}}$; cf. the term $(q+1)\log(q/\delta)$ in [Hin92, line (1.24)], where $q=N$ signifies the number of targets $a_{i}$. But fortuitously we were able to modify their proofs by making an additional use of the key pivot relation (6.1.2) particular to our situation of $\{a_{1},\ldots,a_{q}\}=\mu_{N}$.
@@
At this point we have established all the pieces for the proof of our main result. By Theorem 5.3.8, assumption (1) in Proposition 3.0.1 is indeed satisfied, with the sharp constant $A:=\zeta(3)/2>0$. By Theorem 6.0.1 with the choices $p(x):=x^{N}$ and $r:=1-BN^{-3}$, assumption (2) in Proposition 3.0.1 is also satisfied. In terms of the rings of modular forms $M_{N}$ and $R_{N}$ at an even Wohlfahrt level $N$ introduced in 4.2.1, the conclusion of Proposition 3.0.1 is thus an inequality $R_{N}:M_{2}]\leq CN^{3}\log{N}$, for some absolute constant $C\in\mathbf{R}$ independent of $N$. At this point Proposition 4.2.5 proves the equality $R_{N}=M_{N}$ for all $N\in\mathbf{N}$, which is the unbounded denominators conjecture.
The proof of Theorem 1.0.1 is thus completed.
@@
Since $f(\tau)$ is a modular form, we are reduced to the situation of a number field $K$ such that $f(\tau)\in O_{K}\llbracket q^{1/N}\rrbracket$. We use $R_{2N}$ to denote the ring of modular functions with coefficients in $K$, bounded denominators, and cusp widths dividing $2N$. We follow the proof of Proposition 3.0.1 now on the case of the ring $\mathcal{H}(U,x(t),O_{K})$ from Definition 2.0.1: the ring of formal power series $f(x)\in K\llbracket x\rrbracket$ such that $f(x(t))\in O_{K}\llbracket t\rrbracket$ and $L(f)=0$ for some nonzero linear differential operator $L$ over $\mathbf{Q}(x)$ without any singularities on $U$. Then $R_{2N}\subset\mathcal{H}(U,x(t),O_{K})\otimes_{O_{K}}K$. Note that $U$ is stable under the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$, and thus $\mathcal{H}(U,x(t),O_{K})=\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}O_{K}$ and $\dim_{K(p(x))}\mathcal{H}(U,x(t),O_{K})\otimes K(p(x))=\dim_{\mathbf{Q}(p(x))}\mathcal{H}(U,x(t),\mathbf{Z})\otimes\mathbf{Q}(p(x))$. Thus by Theorem 2.0.2, we still have that $R_{2N}$ has dimension at most $CN^{3}\log N$ over $K[\lambda^{\pm},(1-\lambda)^{\pm}]$. The claimed extension to $\overline{\mathbf{Z}}\llbracket q^{1/N}\rrbracket$ Fourier expansions now follows upon remarking that the proof of Proposition 4.2.5 still persists when $\mathbf{Q}$ is replaced by $K$.

### v2

@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
For $N\gg 1$ and $r\in(0,1)$, we have
$\sup_{|x|=r}\log|F_{N}(x)|\ll\frac{N}{1-r},$
where the implicit constants are both absolute.
@@
Set $S(M,N):=\{x\in D(0,1)\,:\,|F_{N}^{N}(x)-1|<e^{-MN}\}$ and set $M=\frac{N}{1-r}$. Since $M\geq 2$, we may take $M_{0}=2$ and $\epsilon=1/2$ in Lemma 5.2.11 and conclude that for $N\gg 1$ (with absolute constant here), we have
$S(M,N)\subset\bigcup_{\gamma\in\Phi_{N}}\gamma\mathcal{H}_{D},\quad\textrm{where }D=\frac{N^{2}M}{\pi^{2}}.$
By Shimizu’s Lemma (see, for example, [EGM98, Theorem 3.1]) and Proposition 5.2.1, we have  $2|c|\cot(\pi/2N)\geq 1$ for all $\widetilde{\gamma}=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\widetilde{\Gamma}_{N}$. Thus by Lemma 5.2.15,
@@
where the implicit constant is absolute. Thus we have $E(\gamma,D)\leq 1-r$ for all $\gamma\in\Phi_{N}$ once $N^{-1}$ times the implicit constant is less than $1$.
By Lemma 5.2.5, the set $\{x\in D(0,1)\,:\,|F_{N}(x)|>e^{M}+1\}$ is contained in $sS(M,N)$, which is contained in $D(0,1)\smallsetminus\overline{D(0,r)}$ by the above argument for $N\gg 1$. Thus we conclude that
$\sup_{|x|=r}\log|F_{N}(x)|\leq\log(e^{M}+1)\ll M=\frac{N}{1-r}.\qed$
@@ 6. Nevanlinna theory and uniform mean growth near the boundary
For our application of Corollary 2.0.5, we prove in this section the following uniform growth bound. Throughout this section, we assume as we may that $N\geq 2$. Then the analytic map $F_{N}:D(0,1)\to\mathbf{P}^{1}$ omits the $N+1\geq 3$ values $\mu_{N}\cup\{\infty\}$. In such a situation, we seek to exploit whatever growth constraints are imposed on the map by Nevanlinna’s value distribution theory. A theorem of Tsuji [Tsu52, Theorem 11] gives the general asymptotic
@@
for any universal covering map $F:D(0,1)\to\mathbf{C}\smallsetminus\{a_{1},\ldots,a_{N}\}$ based at $F(0)=0$ (see also the discussion in Nevanlinna [Nev70, page 272]), however this is only asymptotically in $r\to 1^{-}$ for given punctures $\{a_{i}\}$ whereas we need a uniformity in both $r$ and $N$. It is at the point (6.2.3) of the explicit partial fraction coefficients that our argument below makes a critical use of the special feature of the target set $\mu_{N}\cup\{\infty\}$ of omitted values.
@@
6.1. Preliminaries in Nevanlinna theory
This section collects some standard material from Nevanlinna’s value distribution theory. The reader should feel encouraged to skip this part on a first reading, and refer back as necessary.
6.1.1. The Nevanlinna characteristic
The left-hand side of (6.0.2) is known as the mean proximity function at $\infty$
@@ 6.1.1. The Nevanlinna characteristic
The Nevanlinna characteristic function
$T(r,f):=m(r,f)+N(r,f)$
is the well-behaved quantity functorially.
@@
The Nevanlinna characteristic function $T(r,f)$ satisfies for every $a\in\mathbf{C}$ the relation
@@
6.1.2. The lemma on the logarithmic derivative
The lemma on the logarithmic derivative—a strong explicit form of which is cited in (6.1.12) below—is the centerpiece of R. Nevanlinna’s original analytic proof of his second main theorem of value distribution theory. The logarithmic error feature of this sharp upper bound on the proximity function of a logarithmic derivative enables us to derive Theorem 6.0.1 from the relatively crude supremum growth bound in Lemma 5.2.17.
The reader willing to take (6.1.12) for granted may at this point proceed directly to §6.2. Nevertheless, since the proof simplifies considerably in the case that we need of a functional unit (a nowhere vanishing holomorphic function), we include our own self-contained treatment of a basic explicit case of the lemma on the logarithmic derivative.
@@ 6.1.2. The lemma on the logarithmic derivative
The case of arbitrary meromorphic functions $g:\overline{D(0,R)}\to\mathbf{P}^{1}$ is handled similarly by a differentiation in the general Poisson–Jensen formula, but with rather more work to estimate the finite sum over the zeros and poles of $g$. See for instance [Nev70, §IX.3.1, page 244, (3.2)] or [Hay64, Lemma 2.3 on page 36] for similar bounds. By using a technique due to Kolokolnikov for handling the sum over the zeros and poles, Goldberg and Grinshtein [GG76] obtained the general bound
(6.1.12) $m\Big(r,\frac{g^{\prime}}{g}\Big)<\log^{+}{\Big\{\frac{T(R,g)}{r}\frac{R}{R-r}\Big\}}+5.8501,\quad\textrm{for }g(0)=1,$
@@
For $f:D(0,1)\to\mathbf{C}$ holomorphic, the polar divisor is empty, and so $N(r,f)=0$ and $m(r,f)=T(r,f)$. Since by definition $F_{N}^{N}-1$ is a unit in the ring of holomorphic functions on $D(0,1)$, our requisite bound (6.0.2) rewrites in Nevanlinna notation into
(6.2.1) $T(r,p\circ F_{N})\ll\log{\frac{N}{1-r}},\quad\textrm{for each of }p(x)\in\big\{x^{N}/(x^{N}-1),\,1/(x^{N}-1),\,x^{N}\big\}.$
6.2.1. Equivalence of bounds for different $p(x)$
By Lemma 6.1.1, the fact $x^{N}/(x^{N}-1)=1+1/(x^{N}-1)$, and (6.1.4), the three cases for $p(x)$ are equivalent to one another. Here we give the explicit estimate in one direction, which will be used later:
$=T\Big(r,\frac{1}{F_{N}^{N}-1}\Big)-\log{2}=T(r,F_{N}^{N}-1)-\log{2}$
(6.2.2) $\geq T(r,F_{N}^{N})-2\log{2}=N\,T(r,F_{N})-\log{4},$
where we use $F_{N}(0)=0$ and $F_{N}^{N}-1$ is a unit in the ring of holomorphic functions on $D(0,1)$. In the rest of this subsection, we will prove Theorem 6.0.1 in the form $T(r,F_{N}^{N})\ll\log{\frac{N}{1-r}}$ but pivoting around the choice
6.2.2. Reduction to a logarithmic derivative
(6.2.4) $f:=1-F_{N}^{N}\quad:\quad D(0,1)\to\mathbf{C}^{\times}$
is related to $p\circ F_{N}=F_{N}^{N}/(F_{N}^{N}-1)$ by
(6.2.5) $p\circ F_{N}=\frac{F_{N}}{NF_{N}^{\prime}}\frac{f^{\prime}}{f}.$
The idea then is that the piece $f^{\prime}/f$ in the decomposition 6.2.5 is small on average over circles by the lemma on the logarithmic derivative (Corollary 6.2.6 below), while—again by the lemma on the logarithmic derivative, in Corollary 6.2.8 below—the characteristic functions of $p\circ F_{N}$ and $F_{N}/F_{N}^{\prime}$ are equal respectively to $NT(r,F_{N})$ and $T(r,F_{N})$ up to a small error.
6.2.3. Two corollaries of the lemma on the logarithmic derivative
Corollary 6.2.6.
For $f=1-F_{N}^{N}$, we have
(6.2.7) $m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log^{+}{\frac{N}{1-r}},$
Proof.
By applying Lemma 6.1.5 to $f$ and the outer radius choice R:=1(1r)/2=(1+r)/2R:=1-(1-r)/2=(1+r)/2, and using (cf. [BG06, Corollary 13.2.14]) that $m(r,f^{\prime}/f)=T(r,f^{\prime}/f)$ is a monotone increasing function of $r$, we find the mean growth bound
@@ 6.2.3. Two corollaries of the lemma on the logarithmic derivative
$\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log^{+}{\frac{N}{1-r}},$
where in the last step we have estimated a mean proximity function trivially by a supremum function. ∎
Corollary 6.2.8.
We have
(6.2.9) $m\Big(r,\frac{F_{N}}{F_{N}^{\prime}}\Big)\leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big).$
The idea of the proof is to combine Lemma 6.1.5 applied to the functional unit $1-F_{N}$ and the standard chain of implications based on Jensen’s formula in the reduction of the second main theorem to the lemma on the logarithmic derivative (see, for example, [Hay64, pages 33–34]).
Proof.
By (6.1.3) for the function $F_{N}^{\prime}/F_{N}$, and the fact that $F_{N}$ is holomorphic on the disc $D(0,1)$ with $F_{N}(0)=0$ and $F_{N}^{\prime}(0)\neq 0$, we have:
@@
(6.2.10) $=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big)-N\Big(r,1/F_{N}^{\prime}\Big)=m\Big(r,\frac{F_{N}^{\prime}}{F_{N}}\Big)+N\Big(r,1/F_{N}\Big).$
Here for the last equality we recall that $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ is an étale analytic mapping, hence the derivative $F_{N}^{\prime}$ is nowhere vanishing.
@@
$\leq T(r,F_{N})+O\Big(\log^{+}{\frac{N}{1-r}}+\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}\Big),$
upon again using Lemma 6.1.5 with R:=(1r)/2R:=(1-r)/2 but now for the functional unit $g=1-F_{N}$, and a similar argument as in the proof of Corollary 6.2.6. ∎
6.2.4. Completing the proof from the crude supremum bound in Lemma 5.2.17
At this point the key identity (6.2.5) allows to combine the estimates (6.2.7) and (6.2.9), arriving at the uniform bound
@@ 6.2.5. A historical note
outside of an exceptional set of radii $r\in E\subset[0,\infty)$ of finite Lebesgue measure: $m(E)<\infty$. Here $N_{\mathrm{ram}}(r,g)=N(r,1/g^{\prime})$ is a ramification term, which is always nonnegative and vanishes if the map $g$ is étale. This is Nevanlinna’s quantitative strengthening of Picard’s theorem on at most one omitted value for a nonconstant entire function, for if each of $a_{1},\ldots,a_{N}$ is omitted then all counting terms $N(r,a_{i})=0$ vanish on the right-hand side of (6.2.13), leading if $N\geq 2$ to an $O(\log{r})$ upper bound on the growth $T(r,g)$ of $g$. The idea is that we similarly have a holomorphic map $F_{N}$ omitting the $N$ values $a_{h}=\exp(2\pi ih/N)$, except $F_{N}$ is on a disc rather than the entire plane, and that (6.2.13) largely extends as a growth bound for holomorphic maps on a disc. For such completely quantitative results we refer the reader to Hinkkanen [Hin92, Theorem 3] or Cherry–Ye [CY01, Theorem 4.2.1 or Theorem 2.8.6]. We cannot directly apply these general theorems in their verbatim forms as they only lead to a bound of the form $m(r,F_{N})\ll\frac{1}{N}\log{\frac{1}{1-r}}+\log{N}$ in place of the required $m(r,F_{N})\ll\frac{1}{N}\log{\frac{N}{1-r}}$; cf. the term $(q+1)\log(q/\delta)$ in [Hin92, line (1.24)], where $q=N$ signifies the number of targets $a_{i}$. But fortuitously we were able to modify their proofs by making an additional use of the key pivot relation (6.2.3) particular to our situation of $\{a_{1},\ldots,a_{q}\}=\mu_{N}$.
For our case of functions on the disc, we compare to  [Hay64, Theorem 2.1]. For holomorphic $f$, we again have the ramification term $N_{\mathrm{ram}}(r,f)=N(r,1/f^{\prime})$ (this term is denoted by $N_{1}(r)$ in [Hay64]), which is always nonnegative. In (6.2.10), even without using the étaleness of $F_{N}$, one would drop the ramification term by positivity and still obtain the requisite bound $m\Big(r,\frac{\varphi}{\varphi^{\prime}}\Big)\leq m\Big(r,\frac{\varphi^{\prime}}{\varphi}\Big)+N\Big(r,1/\varphi\Big)$. In this way, our treatment also recovers the bound $\int_{|z|=r}\log^{+}{|\varphi|}\,\mu_{\mathrm{Haar}}\leq\frac{1}{N-1}\log{\frac{1}{1-r}}+O_{\varphi}(1)$ for every holomorphic map $\varphi:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ avoiding the $N$-th roots of unity (which is not necessarily the universal covering map).
@@
At this point we have established all the pieces for the proof of our main result. By Theorem 5.1.4, assumption (1) in Proposition 3.0.1 is indeed satisfied, with the sharp constant $A:=\zeta(3)/2>0$. By Theorem 6.0.1 with the choices $p(x):=x^{N}$ and $r:=1-BN^{-3}$, assumption (2) in Proposition 3.0.1 is also satisfied. In terms of the algebras of modular forms $M_{N}$ and $R_{N}$ at an even Wohlfahrt level $N$ introduced in 4.2.1, the conclusion of Proposition 3.0.1 is thus an inequality $R_{N}:M_{2}]\leq CN^{3}\log{N}$, for some absolute constant $C\in\mathbf{R}$ independent of $N$. At this point Proposition 4.2.7 proves the equality $R_{N}=M_{N}$ for all $N\in\mathbf{N}$, which is the unbounded denominators conjecture.
The proof of Theorem 1.0.1 is thus completed. ∎
@@
Since $f(\tau)$ is a modular form, we are reduced to the situation of a number field $K$ such that $f(\tau)\in O_{K}\llbracket q^{1/N}\rrbracket$. We use $R_{2N}$ to denote the $K(\lambda)$-algebra generated by modular functions with coefficients in $K$, bounded denominators at $\zeta=i\infty$, and cusp widths dividing $2N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$ (similar to Definition 4.2.1). We follow the proof of Proposition 3.0.1 now on the case of the $K(\lambda)$-vector space $\mathcal{V}(U,x(t),O_{K})$ from Definition 2.0.4. Then $R_{2N}\subset\mathcal{V}(U,x(t),O_{K})$. Note that $U$ is stable under the action of $\mathrm{Gal}(\overline{\mathbf{Q}}/\mathbf{Q})$, and thus $\mathcal{V}(U,x(t),O_{K})=\mathcal{V}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Q}}K$ and $\dim_{K(\lambda)}\mathcal{V}(U,x(t),O_{K})=\dim_{\mathbf{Q}(\lambda)}\mathcal{V}(U,x(t),\mathbf{Z})$. Thus by Corollary 2.0.5, we still have that $R_{2N}$ has dimension at most $CN^{3}\log N$ over $K(\lambda)$. The claimed extension to $\overline{\mathbf{Z}}\llbracket q^{1/N}\rrbracket$ Fourier expansions now follows upon remarking that the proof of Proposition 4.2.7 still persists when $\mathbf{Q}$ is replaced by $K$.
Remark 6.3.2.
It is also possible to derive the $\bar{\mathbf{Z}}\llbracket q^{1/N}\rrbracket$ generalization directly from Theorem 1.0.1, by the following argument pointed out to us by John Voight. The absolute Galois group $\mathrm{Gal}(\bar{\mathbf{Q}}/\mathbf{Q})$ acts on the $q$-expansions of modular forms. If $f(\tau)\in O_{K}\llbracket q^{1/N}\rrbracket$ is a modular form on a finite index subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$, and $\alpha_{1},\ldots,\alpha_{d}$ is a $\mathbf{Z}$-basis of $O_{K}$, then $f_{i}(\tau):=\mathrm{Tr}_{K/\mathbf{Q}}(\alpha_{i}f(\tau))\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ for each $i=1,\ldots,d$. Theorem 1.0.1 gives that each $f_{i}(\tau)$ is modular for some congruence subgroup, say $\Gamma_{i}$. At this point $f(\tau)$, being a $K$-linear combination of $f_{1},\ldots,f_{d}$, is modular for the congruence subgroup $\Gamma_{1}\cap\cdots\cap\Gamma_{d}$.